OpenStudy (anonymous):

Short proof that sin(x)=sin(180-x)?

hartnn (hartnn):

you know about the \(\Large \sin (A+B)\) formula ?

hartnn (hartnn):

** \(\sin (A-B)\)

OpenStudy (anonymous):

I don't wanna remember too many identities. I just don't want to forget the sin(x)=sin(180-x). How can I see it graphicly in cartesian coordinates? :(

hartnn (hartnn):